A firm has a bond issue outstanding with 15 years to maturity and a coupon rate of 8 percent (payments are semiannual). The par value of each bond is $1,000. The required rate has now risen to 12 percent per year. What is the current value of each bond?
Answer #1

Coupon = (0.08 * 1000) / 2 = 40

Rate = 12% / 2 = 6%

Number of periods = 15 * 2 = 30

current value of each bond = Annuity * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

current value of each bond = 40 * [1 - 1 / (1 + 0.06)30] / 0.06 + 1000 / (1 + 0.06)30

current value of each bond = 40 * [1 - 0.17411] / 0.06 + 174.110131

current value of each bond = 40 * 13.764833 + 174.110131

current value of each bond = $724.70